# DeciMamba: Exploring the Length Extrapolation Potential of Mamba

### v1

@@ 6 Conclusions
Looking ahead, we plan to explore different transformer context-extension methods, including length-extrapolation PE Press et al. (2021); Golovneva et al. (2024), hierarchical models, and architectural improvements Sun et al. (2022). Furthermore, we intend to extend our analysis to other layers beyond S6, e.g., RWKV, xLSTM, and more.
@@ 7 Limitations
Our model modifies a pretrained Mamba model but does not propose an improved Mamba architecture to address the underlying issue. Despite empirical evidence demonstrating the effectiveness of our decimation-based method in capturing long-range interactions and its efficiency, the approach can be suboptimal. For example, pooling and compression methods may miss critical information in challenging scenarios. Therefore, designing an improved Mamba variant with enhanced length-generalization abilities that can effectively capture global interactions within a single layer (without pooling the sequence) is an important next step.
@@ 8 Ethics Statement
This work analyzes and improves the performance of large language models (LLMs) for long context understanding, which is crucial when deploying them in real-world systems. This improvement is anticipated to have a positive impact on the use of LLMs in society. However, we acknowledge that LLMs could propagate biases. We emphasize the necessity of further research into these biases before our work can be applied reliably beyond the research environment.
Acknowledgments
This work was supported by a grant from the Tel Aviv University Center for AI and Data Science (TAD), and was partially supported by the KLA foundation. We would like to thank Yael Vinker and Maor Ivgi for fruitful discussions and valuable input which helped improve this work.
@@ A.1 Passkey Retrieval
Each model is trained for 5 epochs with a learning rate of 1e-4, gradient clipping of 1, batch size of 32 (used batch accumulation) and AdamW optimizer Kingma and Ba (2017) with weight decay of 0.1. In each epoch the models train over 6144 sequences of length 2K. For DeciMamba-130M we use $L\_base=2K$, $\beta=0.5$, $decimating\_layers=[13,...,21],min\_seq\_len=20$.
@@ A.2 Document Retrieval
We train each model with data from SQuAD v2 Rajpurkar et al. (2018), which provides examples in the form of (Query, Document, Answer). Our training samples have the following form: <$Query$>; $N_{docs}\times$<$Document$>; <$Answer$>, where <$Document$> can be either the golden document (which holds the answer to the query) or one of $N_{docs}-1$ randomly sampled documents. <$Answer$> holds the id of the golden document. In our setting $N_{docs}=11$, the order of the documents is random, and the query and respective document id are appended to the beginning of each document. During Evaluation we use the same setting but vary the value of $N_{docs}$, between 11 and 300. We note that an average document in SQuAD has a length of about 200 tokens, so our average training sample has about 2,000 tokens, and the evaluation samples vary between 2,000 tokens to 60,000 tokens. We train for one epoch with 300 steps, use a learning rate of 1e-4, gradient clipping of 1, batch size of 64 (used batch accumulation) and AdamW optimizer with weight decay of 0.1.
@@
A.3 Multi-Document Question Answering With Free Text Response
We operate in a similar setting as in Section A.2, but instead of predicting the tokens of the id of the relevant document we let the model generate a free-text response and measure it’s F1 score w.r.t a set of ground truth answers. We train each model for one epoch on the full SQuAD train set (about 90,000 examples when leaving out the samples intended for negative sampling, which do not have a ground-truth answer). We found that the optimal decimation parameters are decimation_layer = 14, $L_{base}$=2000 during training and $L_{base}$=7000 during evaluation. We intentionally decreased Lbase during training so the model could experience decimation during the training period ($L_{train}$ was a bit higher than $L_{base}$), because otherwise the training of DeciMamba and Mamba would have been identical. We use a learning rate of 1e-4, gradient clipping of 1, batch size of 64 (used batch accumulation) and AdamW optimizer with weight decay of 0.1.
A.4 PG-19 Perplexity
We train each model on a total of 100M tokens with a learning rate of 1e-4, gradient clipping of 1, batch size of 250 (used batch accumulation) and AdamW optimizer with weight decay of 0.1. During training we sample a single window from each example and train on it (For the extrapolating models the window length is 2K, for the lower bound models the window length is equal to the context length trained on). During evaluation, for each example we evaluate 10 windows with a maximal constant stride. We evaluate only the last 100 labels in each window, which represent the extrapolation abilities of the model at sequence lengths in the range of [$ctx\_len-100,ctx\_len$], providing an approximation to the model’s performance at the wanted $ctx\_len$. For DeciMamba-130M we use $L\_base=2K$, $\beta=0.83$, $decimating\_layers=[12,...,20],min\_seq\_len=20$. During evaluation we keep the same parameters except setting $L\_base=8K$. Additionally, in this specific task DeciMamba was trained with a similar, yet not identical, Language Modeling (LM) loss. We break the labels sequence (length = 2K) into two chunks. The first 1K labels are trained conditionally on the first 1K tokens of the sequence (like vanilla LM). The last 1K labels are trained conditionally on the whole sequence (2K), and DeciMamba was configured to compress the first 1K input tokens. This way we are able to train DeciMamba to compress context while training on each label in the sequence, making the training much more efficient. We also experimented with chunking the labels into more than two chunks, but only experienced a slowdown in computation while achieving similar performance. For the lower bound models we had to reduce the amount of training steps in order to constrain the training to 100M tokens. Specifically, for each context length, we followed the following formula: $num\_of\_steps=100M/(batch\_size*ctx\_len)=100M/(250*ctx\_len)$. For the Zero-Shot perplexity test for the 1.4b model we used Layer 12 for decimation and $L_{base}=4000$. For the 2.8b model we used Layer 22 for decimation and $L_{base}=4000$.

### v2

@@ 6 Conclusions
Looking ahead, we plan to explore different transformer context-extension methods, including length-extrapolation PE (press2021train; golovneva2024contextual), hierarchical models, and architectural improvements (sun2022length). Furthermore, while the techniques described in this paper are designed for Mamba models, a detailed discussion on extending our proposed methods to other models such as Griffin can be found in Appendix F. Finally, we will further use our Mamba Mean Distance Metric to investigate the mechanism which leads to limited ERFs in Mamba.
@@ 7 Limitations
Similar to other context-extension methods, our model modifies a pretrained Mamba model but does not propose an improved Mamba architecture to address the underlying issue. Despite empirical evidence demonstrating the effectiveness of our decimation-based method in capturing long-range interactions and its efficiency, the approach can be suboptimal. For example, pooling and compression methods may miss critical information in challenging scenarios. Therefore, designing an improved Mamba variant with enhanced length-generalization abilities that can effectively capture global interactions within a single layer (without pooling the sequence) is an important next step.
8 Reproduceability Statement
First, we provide the source code used for the key experiments. Second, in appendix A we provide full configurations for all experiments including instructions on how to train and evaluate the models. We also explicitly state all used datasets, models and hardware, and describe the data processing pipeline and exact metrics used in each experiment.
@@ 9 Ethics Statement
This work analyzes and improves long-context understanding, which is crucial when deploying LLMs in real-world systems. This improvement is anticipated to have a positive impact on the use of LLMs in society. However, we acknowledge that LLMs can propagate biases. We emphasize the necessity of further research into these biases before our work can be applied reliably beyond the research environment.
@@ A.1 Passkey Retrieval
Each model is trained for 5 epochs with a learning rate of 1e-4, gradient clipping of 1, batch size of 32 (used batch accumulation) and AdamW optimizer (kingma2017adam) with weight decay of 0.1. In each epoch the models train over 6144 sequences of length 2K. For DeciMamba-130M we use $L\_base=2K$, $\beta=0.5$, $decimating\_layers=[13,\dots,21],min\_seq\_len=20$. Our code is built over an existing version of BABILong (kuratov2024babilong).
@@ A.2 Document Retrieval
We train each model with data from SQuAD v2 (rajpurkar2018squadv2), which provides examples in the form of (Query, Document, Answer). Our training samples have the following form: $N_{docs}\times$$<$$Document$$>$; $<$$Answer$$>$, where $<$$Document$$>$ can be either the golden document (which holds the answer to the query) or one of $N_{docs}-1$ randomly sampled documents. $<$$Answer$$>$ holds the id of the golden document. In our setting $N_{docs}=12$, the order of the documents is random, and the query and respective document id are appended to the beginning of each document. During Evaluation we use the same setting but vary the value of $N_{docs}$, between 12 and 300. We note that the length of an average document in SQuAD is a bit smaller than 200 tokens, so our average training sample has about 2,200 tokens, and the evaluation samples vary between 2,200 tokens to 60,000 tokens. We train for one epoch with 300 steps, use a learning rate of 1e-4, gradient clipping of 1, batch size of 64 (used batch accumulation) and AdamW optimizer with weight decay of 0.1. We found that the optimal decimation parameters are decimation_layer = 12, $L_{base}$=2000 during training and $L_{base}$=4000 during evaluation. We intentionally decreased Lbase during training so the model could experience decimation during the training period ($L_{train}$ was a bit higher than $L_{base}$), because otherwise the training of DeciMamba and Mamba would have been identical.
A.3 LongBench
We use the ‘iuyul/mamba-2.8b-zephyr’ instruction-tuned model from huggingface. decimation_layer = 28. $L_{base}$ varies between different benchmarks, between 2000 to 2800.
A.4 PG-19 Perplexity
We train each model on a total of 100M tokens with a learning rate of 1e-4, gradient clipping of 1, batch size of 250 (used batch accumulation) and AdamW optimizer with weight decay of 0.1. During training we sample a single window from each example and train on it (For the extrapolating models the window length is 2K, for the lower bound models the window length is equal to the context length trained on). During evaluation, for each example we evaluate 10 windows with a maximal constant stride. We evaluate only the last 100 labels in each window, which represent the extrapolation abilities of the model at sequence lengths in the range of [$ctx\_len-100,ctx\_len$], providing an approximation to the model’s performance at the wanted $ctx\_len$. For DeciMamba-130M we use $L\_base=2K$, $\beta=0.83$, $decimating\_layers=[12,\dots,20],min\_seq\_len=20$. During evaluation we keep the same parameters except setting $L\_base=8K$. Additionally, in this specific task DeciMamba was trained with a similar, yet not identical, Language Modeling (LM) loss. We break the labels sequence (length = 2K) into two chunks. The first 1K labels are trained conditionally on the first 1K tokens of the sequence (like vanilla LM). The last 1K labels are trained conditionally on the whole sequence (2K), and DeciMamba was configured to compress the first 1K input tokens. This way we are able to train DeciMamba to compress context while training on each label in the sequence, making the training much more efficient. We also experimented with chunking the labels into more than two chunks, but only experienced a slowdown in computation while achieving similar performance. For the lower bound models we had to reduce the amount of training steps in order to constrain the training to 100M tokens. Specifically, for each context length, we followed the following formula: $num\_of\_steps=100M/(batch\_size*ctx\_len)=100M/(250*ctx\_len)$. For the 1.4b model we used Layer 12 for decimation and $L_{base}=4000$. For the 2.8b model we used Layer 22 for decimation and $L_{base}=4000$.
Appendix B Additional Evaluations
B.1 Comparison With Transformers
We find that vanilla Transformers of equivalent size (trained on the same dataset with a similar training recipe), have inferior length generalization abilities compared to Mamba. This is evident in all long-context tasks tested (Tables 4, 5, 6).
Table 4: Comparison With Transformers - Passkey Retrieval. The setting is the same as in AppendixA.1. All models were trained on sequences of length 2k. For each context length we test performance for 5 different needle locations, and report the success rate (between 0 and 1). ✓stands for 100% success, ✗ for 0% success.
Context Length 1K 2K 4K 8K 16K 32K 64K 128K
Pythia-160M
Mamba-130M 0.8
DeciMamba-130M
Table 5: Comparison With Transformers - Multi-Document Retrieval. The setting is the same as in AppendixA.2. All models were trained on sequences of length 2k (10 documents). For each context length we test performance for 100 different queries and report the success rate (between 0 and 1).
Num of Docs 10 20 40 60 80 100 120 140 160 180 200 250 300
Pythia-160M 0.69
Mamba-130M 0.59 0.61 0.41 0.06
DeciMamba-130M 0.70 0.70 0.53 0.48 0.51 0.58 0.67 0.68 0.73 0.60 0.61 0.53 0.21
Table 6: Comparison With Transformers - Zero-Shot Perplexity. The setting is the same as in Appendix A.4. All models were pre-trained on sequences of length 2k (we remind that DeciMamba is applied directly without any tuning). inf replaces any perplexity result larger than 100.
Context Length 1K 2K 4K 8K 10K 20K 30K 40K 50K 60K 70K 80K
Pythia-2.8B 10.24 9.96 inf inf inf inf inf inf inf inf inf inf
Mamba-2.8B 9.39 9.17 11.6 inf inf inf inf inf inf inf inf inf
DeciMamba-2.8B 9.39 9.17 11.98 14.58 14.73 17.17 19.83 22.2 24.89 27.57 27.89 27.43
Pythia-1.4B 11.64 11.32 inf inf inf inf inf inf inf inf inf inf
Mamba-1.4B 10.51 10.31 10.5 14.43 18.4 inf inf inf inf inf inf inf
DeciMamba-1.4B 10.51 10.31 10.5 14.13 14.5 18.33 23.54 26.82 28.97 30.56 29.97 29.28
For LongBench, the equivalent Transformer model (Pythia-2.8B) repeatedly causes an OOM error on our GPU (A6000, 48GB of RAM).
Appendix C Hyperparameter Selection
In practice, $L_{base}$ is the only parameter we sweep. We typically scan 3 or 4 values that are similar in magnitude to the context length used during training ($L_{train}$). The reason for selecting $L_{base}$ close to $L_{train}$ originates from an assumption on the long-context data: short training sequences and long evaluation sequences have similar information content and differ mainly by the amount of noise in the sequence. Under this assumption, there are at most $L_{train}$ important tokens; hence, it makes sense to pool this number of tokens, regardless of the amount of noise. We find this assumption quite reasonable for many long-context tasks, such as retrieval, multi-document question answering, and next-token prediction. For example, in multi-document retrieval/QA, only one document is relevant to the query, regardless of how many random documents we append to the context. Another example is next-token prediction, which is usually very local and does not benefit much from global interactions. Another reason for selecting $L_{base}$ close to $L_{train}$ is that there are no ERF issues when processing sequences of length $L_{train}$, as the global layers (starting from the first decimation layer) are trained on sequences of the same length.
Regarding the number of layers to decimate and the decay rate ($\beta$): The main goal of these parameters is to improve efficiency, not performance. Both parameters introduce the option to compress the sequence further by applying additional DeciMamba layers. We emphasize that this has negligible effects on performance, as shown in the sweeps below for the passkey retrieval task (Tables 7, 8). In addition, this is supported by other results, such as Multi-Document Retrieval and LongBench, which were run with this option disabled.
Table 7: Decay Rate ($\beta$) Sweep. BL is the performance of the baseline Mamba model. The setting is the same as in AppendixA.1. We use $L_{base}=2000$ and 9 decimation layers (13-21).
$\beta$ BL 1 0.75 0.5 0.25
Success Rate 0.55 1.0 0.975 0.975 0.975
Table 8: Number of Decimation Layers Sweep. BL is the performance of the baseline Mamba model. The setting is the same as in Appendix A.1. We use $L_{base}=2000$ and $\beta=0.5$
# of Deci Layers BL 1 2 3 4 5 6 7 8 9
Success Rate 0.55 1.0 0.975 0.925 0.925 0.975 0.925 0.95 0.95 0.975
Appendix D Other Related Work
D.1 Long Range Transformers.
Transformers have emerged as highly effective models for various tasks, yet their widespread adoption has been constrained by their limited long-range modeling capabilities. Thus, applying transformers effectively to long-range data remains a central challenge in DL, particularly in NLP. A primary factor in this challenge is that the effective context of transformers is dominated by the context observed during training, which is limited because training LLMs on datasets with billions of tokens across lengthy sequences is computationally demanding. Hence, three main approaches have been developed to tackle this problem: (i) creating efficient variants of transformers that allow an increase in the length of sequences during training. (ii) Context extension methods, which enable training on short sequences and evaluation on long sequences, and finally, (iii) hierarchical models that rely on pooling, chunking, and compression. Despite these extensive efforts, several recent studies indicate that high-quality handling of long text remains an unresolved issue (liu2024lost; li2024long).
D.2 Efficient transformers.
Over the years, many approaches have been proposed for making transformers more efficient (tay2022efficient; fournier2023practical). The two most prominent directions are hardware-aware implementations such as flash-attention (dao2022flashattention; dao2023flashattention) and ring-attention (liu2023ring), which accelerate computations over long sequences by several orders of magnitude. Additionally, developing efficient attention variants with sub-quadratic complexity has become very popular. Two notable examples are Linformer (wang2020linformer), which utilizes a low-rank attention matrix, and Performer (choromanski2020rethinking), a variant that approximates the attention operator through a kernel function.
D.3 Effective Receptive Field (ERF)
The Receptive Field (RF) (lecun1998gradient) of a neuron refers to the region of the input space that influences the neuron’s output. In traditional CNNs, the RF captures how the area of influence expands as the neuron’s depth increases. However, in modern architectures like transformers, which process the entire input context at each layer and can theoretically model interactions between any parts of the input, the concept of RF becomes less informative. In these models, the theoretical receptive field is maximal, rendering the term ambiguous.
To address this, the ERF offers a more practical measure by quantifying the empirical sensitivity of a model to different input regions. The ERF highlights where the model actually focuses its attention and depends not only on the architecture but also on the data and training process. The seminal work on Vision Transformer (ViT) (dosovitskiy2020image) and vig2019analyzing introduced the concept of attention mean distance to measure the ERF, as detailed in Eq. 6. Building upon this, we extend the measurement to Mamba layers in Eq. 8, allowing us to effectively estimate their ERF.
D.4 Why use Long Context LLMs instead of Retrieval Augmented Generation?
Performance-wise, recent studies do not show a conclusive advantage of RAG over Long Context LLMs: li2024retrievalaugmentedgenerationlongcontext, xu2024retrievalmeetslongcontext. Interestingly, the common conclusion of all these studies (and others, such as Anthropic’s Contextual Retrieval) is that when combined, RAG and Long-Context LLMs have a synergistic effect which improves performance even more. Moreover, each study proposes a different way of combining RAG and Long Context LLMs (directly, switching between RAG and Long Context, augmenting the retriever chunks, etc.) - showing that further research in each area alone, or combined, has potential for an overall improvement in LLM performance.
D.5 Justification for Mamba Mean Distance
We are aware of two main approaches for measuring the ERF of Mamba and Attention models: (i) via the attention mean distance or our analogy of Mamba Distance (Eq. 10), and (ii) by measuring the gradient norms over the entire model, which was employed for transformers in chi2022dissecting and very recently applied to Mamba models (jafari2024mambalrp) (section 6). The main drawback of the second approach is that it measures the ERF in the context of gradients, which are significantly influenced by the type of loss function, task, and label, complicating the analysis. Additionally, while the first method can be applied to a single layer or a set of layers, the second method becomes less straightforward when applied to a subset of layers, requiring additional considerations and design choices, such as whether to normalize gradients, the use of gradients instead of relevance scores, and the computation of gradients only for attention blocks that learn token interactions or also consider linear and normalization layers. Due to its flexibility and simplicity, as well as being more popular (several examples are vig2019analyzing; dosovitskiy2020image; raghu2021vision), we chose the first method.
Appendix E Ablations
E.1 Layer Selection
We ablate our layer selection methodology, as it has significant affect on the model’s performance (Figure 8, left). Each point on the curve represents the score achieved by a DeciMamba-130M model in the Passkey Retrieval task when trained on sequences of length 2K. The only difference between the models is the first decimating layer (x axis). We see that when the decimating layer is too shallow (e.g. layer 8), the model fails completely. As we increase the minimal decimating layer we observe a large increase in performance, until reaching the climax at layer 12. This result aligns with the fact that the $\Delta_{t}$ distribution is farther from 0 in Mamba’s early layers (all tokens are important). We hypothesize that the tokens still don’t have a strong representation at this stage, hence it makes less sense to decimate w.r.t their $\Delta_{t}$ value. After layer 12 the performance of DeciMamba starts to drop, and stabilizes at the vicinity of the baseline Mamba model. At this region we start decimating too late, as some long-dependency layers suffer from ERFs when processing the longer Passkey Retrieval sequences. The final layers (21-23) also have larger $\Delta_{t}$ values, so they should not be decimated as well. We hypothesize that these layers ’decode’ the processed embeddings back into token distributions, so the intermediate representations are yet again not fit for semantic decimation.
E.2 Pooling Strategy
We show the importance of the proposed Top-$P_{s}$ pooling strategy in the following ablation study. Suppose we select a different pooling strategy, Top-$K\%$, which keeps the top $K\%$ tokens with the largest mean $\Delta_{t}$ values (over the channels dimension). We train a DeciMamba-130m model for the Passkey Retrieval task on sequences of length $L_{train}=2K$, with $K=50\%$ and decimation layers $[12,\dots,20]$. As can be seen in Fig. 9 (Appendix), Top-$K\%$ extends Mamba’s extrapolation abilities from 8K to 32K, yet does not do as well as Top-$P_{s}$ which extends to 128K. While Top-$K\%$ decimation decreases the sequence length greatly (the output is x512 times shorter), during extrapolation each SSM layer still faces input sequences that are much longer than the ones it has seen during training, as demonstrated in Fig. 10. We further explain this result; During training Layer 17 saw sequences of length 65 ($L_{train}=2K$, followed by a decimation rate of 50% in each layer from layer 12). During inference it sees sequences of lengths 65 and 1002 (where $L_{eval}=2K,32K$, left and right images, respectively). Since the sequence length at layer 17 of the latter (1K) is much longer than the one it saw during training (65), the extrapolation is limited by an ERF (right image, dashed orange shape). This shows the importance of keeping the lengths of the inputs to each decimating SSM layer similar to the ones seen during training, as done by Top-$P_{s}$.
@@ E.2 Pooling Strategy
E.3 Decimation Mechanism.
To emphasize the benefit of pooling w.r.t the values of $\Delta_{t}$, we examine two additional decimation mechanisms: max-norm decimation (tokens with maximal norm are kept) and random decimation (tokens are randomly kept). We equip Mamba-1.4b with each of the decimation mechanisms (decimation layer = 12, $L_{base}=8000$), and compare each model’s zero-shot perplexity in Fig. 8 (right). Max-norm pooling achieves a similar trend to $\Delta_{t}$ pooling (DeciMamba), yet performs somewhat worse. The result suggests that $\Delta_{t}$ is a better candidate for pooling, but also shows that the token’s norm also holds some information about its importance. We hypothesize that a method that combines both might achieve better performance than any one alone, but leave this question open for future work. Random pooling induces strong distortion to the embedded sequence, yet achieves better perplexity for longer contexts. This surprising result demonstrates how sensitive Mamba is to limited ERFs - in longer sequences, it is actually better to randomly decimate the sequence rather than process its full length.
E.4 Adding Additional Pooling During Decoding.
By design, DeciMamba performs pooling only during the pre-fill phase of inference. Although the approach is motivated by many previous works (jiang2024minference10acceleratingprefilling; ge2024incontextautoencodercontextcompression; jiang2024longllmlinguaacceleratingenhancingllms; li2023compressingcontextenhanceinference; jiang2023llmlinguacompressingpromptsaccelerated; liu2024cachegenkvcachecompression; wingate2022promptcompressioncontrastiveconditioning; mu2024learningcompresspromptsgist; chevalier2023adaptinglanguagemodelscompress; fei2023extendingcontextwindowlarge; wang2024incontextformerlightningfastcompressing), we further motivate it by ablating an additional pooling mechanism during the decoding phase. We tested this on the longest generation task in LongBench: GovReport, which involves summarizing long documents with an average length of 10,000 tokens. The additional decimation during decoding was performed by iteratively combining prefill and decoding for each generated chunk, where the chunk size is 50 tokens. Table 9 presents the results and shows that adding decoding pooling has a negligible effect on performance.
A possible explanation for the limited impact of adding decimation during decoding is that, despite being long (about 500 tokens), the generated summaries are relatively short compared to the number of tokens that cause limited ERFs (it needs to be longer than the context length used to train Mamba). Thus, the above suggests that decimation is not needed during decoding. Note also that while in GovReport the generated responses are relatively long, for most of the remaining 15 long-context tasks in LongBench, such as Multi-Document QA, the model is required to generate only 10 to 30 tokens per answer. Therefore, pooling is not required during their decoding phase as well.
Clearly, if a new task emerges where the number of generated tokens is larger than the ERF, it might be useful to add pooling to the decoding phase.
Table 9: GovReport: Long Summary Generation With and Without Decoding Decimation. LB is LongBench; 0-4k, 4-8k, 8k+ are the three LongBench-e context length groups; and +DD is DeciMamba with additional Decoding Decimation.
Model LB 0-4k 4-8k 8k+
Mamba 9.8 24.8 10.9 5.2
DeciMamba 14.9 27.4 19.4 7.8
DeciMamba + DD 14.7 26.6 18.4 8.14
Appendix F Extensions to Other Models
While this paper focuses on exploring the length-generalization abilities and introducing the first context-extension method specifically designed for Mamba models, our proposed techniques are broadly applicable to other architectures. First, the Mamba Mean Distance metric can be generalized to other models that rely on implicit attention. Examples include HGRN (Eq. 5 in qin2024hierarchically), Griffin (Eq. 14 in de2024griffin), RWKV (Eq. 18 in zimerman2024unified), and others. These tools can provide valuable insights into the ERF of these models, offering a deeper understanding of their length-generalization behavior. Furthermore, the entire context-extension technique introduced in Sec. 4 can be adapted for other models. Specifically, models with per-token data-dependent recurrent gates can leverage these gates to compute token importance scores, enabling a decimation strategy during pre-fill. For instance, in Griffin, the recurrent gate $r_{t}$ (Eq. 1 in  de2024griffin can be used to assign token importance scores, facilitating an effective decimation strategy. Similarly, in HGRN and GLA, the recurrent gate $\lambda_{t}$ (Eq. 2 in qin2024hierarchically) and $G_{t}$ (Eq. 3 in yang2023gated) accordingly can be interpreted as token importance measures, allowing our approach to be extended to design context-extension techniques for these architectures. By extending these techniques, we provide a foundation for exploring length-generalization in various sequence models.
Table 10: Inference Time Per Block. In milliseconds, benchmarked on Nvidia RTX A6000 GPU. DeciMamba and Mamba blocks have a similar processing time. Notice how the processing time drops for blocks that are placed after a DeciMamba block, no matter the input sequence length.
Model Context Length
8K 16K 32K 64K 128K 256K 512K
Mamba Block 3.2 6.1 11.6 22.7 52.5 95.9 204
DeciMamba Block 3.3 6.1 11.8 23.1 45.9 93.1 188
Mamba Following DeciMamba 1.6 1.6 1.5 1.9 1.8 1.9 2.4
DeciMamba Following DeciMamba 1.3 1.4 1.5 1.5 1.5 1.5 1.5